THIS WAS THE ONLY INFORMATION GIVEN, no composition given.. maybe just use 75% though

A 67.2 gg sample of a gold and palladium alloy contains 3.09×1023 atoms.

What is the mass percentage of the gold in the alloy?

What is the mass percentage of the palladium in the alloy?
Answer #1

atomic Let a % of gold = a ditore Arass to of Palladium= 100-x Atomic mass of gold- 196.967 1 Atomic mar af Palladium= 106.42279.7995 n = 40474.56-32883078 x= 90474.56-32883.78 279.7995 = 27.129 Atomic ). of Au= 27.129 % Atonine s. of Pd= (100-27.129